### v2

# Forecasting Arctic Temperatures with Temporally Dependent Data Using Quantile Gradient Boosting and Adaptive Conformal Prediction Regions

Abstract
Using data from the Longyearbyen weather station, quantile gradient boosting (“small AI”) is applied to forecast daily temperatures in Svalbard, Norway. Temperatures above $0^{\circ}\mathrm{C}$ are of special interest because of their impact on ice, snow, and tundra permafrost. To improve forecasting skill for warmer temperatures, the target quantile is 0.60; forecast underestimates are weighted 1.5 times more heavily than forecast overestimates when the quantile loss is computed. Predictors include eight routinely collected indicators of weather conditions, each lagged by 14 days, yielding temperature forecasts with a two-week lead time. Adaptive conformal prediction regions quantify forecasting uncertainty with provably valid coverage. Using a holdout sample, a forecast of $0^{\circ}\mathrm{C}$ is correct 14 days later at least 80% of the time. Implications for Arctic adaptation policy are discussed.
adaptive conformal prediction regions,
keywords:
1 Introduction
The oceans and cryosphere interact to support unique ecosystems while exchanging water, energy, and carbon with Earth’s climate system. The 2019 Intergovernmental Panel on Climate Change (IPCC) report concludes that global warming has altered this interaction, causing “mass loss from ice sheets and glaciers, reductions in snow cover, decreases in Arctic sea-ice extent and thickness, and increased permafrost temperatures” (IPCC, 2019, A.1). Ecosystem impacts have been equally dramatic (IPCC, 2019, A.4).
A key feature of the IPCC report is its reliance on forecasts at very large spatial and temporal scales, consistent with planet-wide coverage and the gradual pace of climate change. Spatial scales can span thousands of square kilometers. [^1] For example, the Community Earth System Model version 2 (CESM2) is typically run at nominal $1^{\circ}$ resolution, corresponding to grid cells of roughly 10,000–15,000 km2, depending on latitude. Downscaling methods have been proposed for regional grids, but they rest on strong assumptions about the parent Global Climate Model (GCM), and no consensus has emerged on a preferred approach (Nishizawa et al., 2018). Moreover, finer spatial grids would greatly increase computational demands that are already near prohibitive. The temporal scale reflects grid cells updated internally every 30 minutes, with atmosphere–land–ocean–ice coupling on hourly intervals and outputs are usually produced as daily or monthly means (Danabasoglu et al., 2020; National Center for Atmospheric Research, 2020). Projections typically extend years into the future.
The goals of numerical weather prediction (NWP) differ from those of climate modeling and provide output at much finer spatial and temporal detail. [^2] Global weather prediction systems, such as ECMWF’s Integrated Forecasting System (IFS) and NOAA’s Global Forecast System (GFS), use grid cells covering on the order of 80–200 km2. Even the most advanced regional models, such as NOAA’s High-Resolution Rapid Refresh (HRRR) or the Weather Research and Forecasting (WRF) model, employ grid spacings of 1–3 km and still aggregate conditions across areas of 1–9 km2. Their time steps are on the order of seconds to minutes, with outputs every 1–3 hours for global models and hourly (or finer) for regional models (European Centre for Medium-Range Weather Forecasts, 2023; NOAA National Centers for Environmental Prediction, 2023; NOAA Earth System Research Laboratory, 2023). Although NWP calculations are performed at much smaller scales, the resulting forecasts still aggregate across those spatial and temporal units. Site-specific forecasts represent averages across larger areas, and hourly predictions are aggregates over seconds and minutes.
The coarse spatial and temporal resolution characterizing GCMs and NWP is especially consequential for forecasts in high-latitude regions such as the Arctic. Small temperature differences around the local freezing point can have important implications, intensified by “Arctic amplification.” Arctic amplification is discussed in more detail shortly.
In addition, the conditions forecasted are usually standardized. The snowfall expected in, say, Fairbanks, Alaska “during the next 24 hours” is one example. Different locations, however, may wish to supplement such forecasts. In the Arctic, local officials might want to know well in advance the probability that $0^{\circ}\mathrm{C}$ will be exceeded. Such information could guide the timing of sediment removal, preemptive slope closures, and preparation of wastewater systems for meltwater discharge, among other actions.
Estimating forecasting uncertainty is problematic as well (Fu, 2025). Gettleman and Rood (2016, 12) summarize the challenges for climate models, which apply equally to weather forecasting: “Uncertainty in climate models has several components. They are related to the model itself, to the initial conditions of the model, and to the inputs that affect the model. All three must be addressed for the model to be useful.”
There have been recent technical advances (Pathak et al., 2022; Price et al., 2024; Bodnar et al., 2025). Yet issues of coarse spatial scales, standardized output, equivocal reliability, and other issues remain largely unresolved (Ortega et al., 2022; Balaji et al., 2022; Chang et al., 2023; Zhang et al., 2025). This paper uses Svalbard, Norway, as a test case to advance Arctic temperature forecasting while addressing these three limitations. Temperatures are the focus because of legitimate concerns about a warming climate.
The statistical methods used here are best viewed as a complement to techniques usually applied at larger spatial and temporal scales. Because these methods are more surgical, they can resolve finer spatial and temporal detail. One can obtain legitimate forecasts coupled with valid measures of forecasting uncertainty well suited for highly local and rapid adaptations to Arctic warming.
Section 2 provides background on Svalbard and the forecasting challenges posed by Arctic amplification. Section 3 describes the data, the construction of holdout samples, and the 14-day lagging of predictors used to support legitimate forecasting. Section 4 introduces the statistical methods, focusing on quantile supervised learning. Section 5 presents the empirical results, emphasizing visualization and algorithm interpretations. Section 6 addresses uncertainty through adaptive conformal prediction regions for multiple time-series data and offers a grounded way to communicate forecast reliability to stakeholders. Section 7 discusses policy implications, with emphasis on short-fuse adaptations to Arctic melting, and Section 8 concludes. A synopsis of the data analysis steps is provided in Appendix A.
2 Svalbard, Norway: The Forecasting Setting
Svalbard is a remote Norwegian archipelago in the Arctic Ocean, located about halfway between mainland Norway and the North Pole. Its main settlement, Longyearbyen, lies at roughly $78^{\circ}\mathrm{N}$ latitude, well above the Arctic Circle. Only a small fraction of the land is vegetated, mostly tundra, while the remainder is dominated by permafrost, ice, and bare rock. Longyearbyen is one of the northernmost permanently inhabited places on Earth.
The maritime climate is distinctive in part because Svalbard is warmed by Atlantic Ocean currents. Average winter temperatures in Longyearbyen range from about $20^{\circ}\mathrm{C}$ to $14^{\circ}\mathrm{C}$, while summers are typically between $3^{\circ}\mathrm{C}$ and $7^{\circ}\mathrm{C}$. In recent years, summer temperatures have more often spiked above $10^{\circ}\mathrm{C}$, a trend climate scientists find troubling.
Snow falls during much of the year, but total precipitation is relatively low – technically making Svalbard a polar desert. From late October to mid-February the archipelago experiences the polar night, when the sun does not rise above the horizon. From mid-April to late August, the sun remains above the horizon, bringing continuous daylight. In recent decades Svalbard has warmed at a rate several times the global average, with mean temperatures increasing by more than $5^{\circ}\mathrm{C}$ since the 1970s. Glaciers are retreating, sea ice is thinning, and thawing permafrost is reshaping the terrain. The archipelago is now among the world’s regions most affected by climate change (Urbański and Litwicka, 2022; Karlsen et al., 2024; Bradley et al., 2025; Schuler et al., 2025).
2.1 Some Implications for Forecasting Temperatures
A practical attraction of Svalbard for temperature forecasting is the weather station at Longyearbyen Airport, whose data are curated by NOAA in collaboration with its international partners. These data are freely available and easy to download. Under the Svalbard Treaty of 1920, Norway retains sovereignty but grants citizens of other signatory nations equal rights to engage in scientific and commercial activities. International research stations are operated by Russia, Poland, Germany, China, the United States, and others.
A scientific attraction is that Svalbard’s maritime climate presents a demanding test for Arctic temperature forecasting. Unlike continental climates at comparable latitudes, one cannot simply look westward and project that those temperatures will arrive several days later. Immediately to the west lies the Arctic Ocean, not a large landmass with comparable terrain.
Another forecasting challenge is the especially rapid climatic change occurring throughout polar regions such as Svalbard, often termed “Arctic amplification” (Rantanen et al., 2022). The farther one moves from the equator, the faster the rate of warming. Early explanations emphasized the declining albedo caused by sea-ice loss and the increasing exposure of tundra and rock. As more sunlight is absorbed and less reflected, additional heat is retained, reinforcing local warming through a strong positive feedback loop.
Semenov (2021) notes that at least two additional mechanisms are now recognized as important contributors to Arctic amplification. Their most fundamental feature is the Planck effect. In simple terms, the Planck effect describes how efficiently Earth radiates heat into space. That efficiency—the rate at which the outgoing energy flux increases with temperature—is smaller in colder regions. Because the radiative flux follows the Stefan–Boltzmann law, its slope with respect to temperature ($T$) is proportional to $T^{3}$. At the low surface temperatures typical of the polar regions, this slope is relatively flat. As a result, the polar surface radiates infrared energy less effectively than warmer regions experiencing the same temperature increase; heat therefore accumulates more rapidly.
The same Planck effect also operates higher in the atmosphere. The “effective emission height” is the average altitude from which the planet’s infrared radiation escapes directly to space. As greenhouse-gas concentrations rise, this emission level shifts upward. Because temperature generally decreases with altitude, radiation then originates from colder air. By the same $T^{3}$ relationship, the efficiency of infrared emission is reduced. Thus, the Planck effect contributes to polar amplification in two ways: directly at the cold surface and indirectly aloft through the upward shift of the emission height to colder temperatures. Arctic amplification provides a potential complication for the forecasting that follows.
3 Data
The data come from the Longyearbyen weather station at the airport in Svalbard, Norway, and are easily downloaded from the Integrated Surface Database using the worldmet package in R. The database contains weather station data from around the globe. It uses the same standard format regardless of origin.
Anticipating the use of adaptive conformal prediction regions (Romano, Patterson and Candès, 2019), a variant on split-sample methods is employed. Observations for 2023 constitute the training data. Calibration data from 2022 provide an “honest” assessment of the number of algorithm iterations required. They are also the observations needed to construct proper nonconformal scores.
The calibration data are divided into observations for which the forecasted temperature is $0^{\circ}\mathrm{C}$ and observations for which the forecasted temperature is $\leq 0^{\circ}\mathrm{C}$. The melting of snow, sea ice, glacier ice, and permafrost creates positive feedback loops that change the manner in which temperature variation is produced. On subject-matter and statistical grounds, adaptive conformal methods are therefore applied separately to each subset. [^3] One might wonder why the data splits are made using forecasted temperatures rather than observed temperatures. Observed temperatures are certainly available in the weather station data. But when forecasting, only the forecast is known. If the future temperature were known, there would be no need to forecast it.
Data from 2024 serve as a pristine holdout sample treated as “new cases” to document true forecasting skill. As test data, these observations have no role whatsoever in training. The data split by forecasted temperature applied to the calibration data is applied again. The requirements necessary to capitalize properly on all three datasets within the split-sample approach are discussed in Section 6. Some data from 2021 are used for lagged-variable construction, briefly explained in footnote 4. They play no other role in the analyses.
The training data, calibration data, and test data from the Integrated Surface Database come with hourly observations within each day. Hourly data are too detailed for the analyses to follow and add unnecessary complexity; daily data are used instead.
The response variable is the daily, solar-time 2 p.m. air temperature in degrees Celsius at the Longyearbyen weather station. Each daily temperature value at 2 p.m. solar time is an instantaneous observation—a snapshot of conditions with a constant daily reference time rather than an hourly or daily mean. The 2 p.m. solar-time convention is used uniformly throughout the year, including during the polar night, to provide a consistent temporal reference for forecasting. This differs from global climate and numerical weather prediction procedures, which typically rely on spatially and temporally averaged fields.
Predictors are all lagged by 14 days because that seems to be a lag commonly used by climate scientists (Li et al., 2024). In future work, longer lags could be considered. The predictors include (1) wind direction in degrees from true north, (2) wind speed in meters per second, (3) air temperature in degrees Celsius, (4) atmospheric pressure in hectopascals (hPa), (5) visibility in meters, (6) dew point in degrees Celsius, (7) relative humidity in percent, and (8) a day counter ranging from 1 to 365. The counter can capture temporal trends: on average, the diurnal months are warmer than the nocturnal months, although temperature changes over time may be nonlinear. Several predictors are likely to interact in complex ways (Semenov, 2021). The 14-day lags translate into fitted 2 p.m. temperatures two weeks later that are a foundation for forecasts 14 days in advance. [^4] Lagging variables is a routine procedure in feature construction. In this case, however, lagging the first 14 days of a year pushes their reference points into the final 14 days of the previous year. Data from 2021 are therefore included to provide lagged values for the first 14 days of 2022. Otherwise, the 2021 data are not used.
The complete dataset forms a multiple time series, making temporal dependence a potential complication. Split samples are commonly disjoint subsets chosen at random from the data available, but that is likely to obscure any temporal dependence (Hyndman and Athanasopoulos, 2021, sec. 5.8). A key requirement for the 2022 and 2024 holdout samples is that the same physical processes apply during the identical months of 2022, 2023, and 2024; data for all three years are treated as random realizations from the same underlying joint probability distribution. Nonetheless, Arctic amplification may affect this comparability, an issue examined empirically in Section 5.2.
4 Statistical Methods
The analyses to follow produce legitimate forecasts from supervised machine learning applied to temporal data. Holdout samples are used to construct provably valid prediction regions. The statistical framework is an interlocking set of procedures. After a search in Google Scholar, their combination appear to be novel. Each component is discussed in a grounded manner when deployed. Appendix A provides a step-by-step synopsis.
4.1 Background
Machine learning algorithms from statistics and computer science are increasingly used in climate and weather applications, including forecasting (Ma et al., 2023; Miloshevich et al., 2024; Kvånum et al., 2025; Zhang et al., 2025). Especially relevant to the analyses that follow is the work of Velthoen et al. (2023), who analyze several years of daily precipitation observations from the Dutch KNMI weather-station network. The precipitation distribution exhibits a long right tail. Weather-station data provide one source of predictors; deterministic precipitation forecasts from the ECMWF weather model provide another. Combined, these data form a multiple time series.
The authors develop gbex, a quantile gradient-boosting algorithm (Friedman, 2001, 2002) tailored for rare and extreme precipitation events using conditional quantiles at very high levels (e.g., 0.95 or 0.99). Quantile regression methods typically struggle when either tail is sparse. The gbex procedure borrows strength from Extreme-Value Theory (EVT) by fitting a Generalized Pareto Distribution (GPD) to values exceeding a high threshold while allowing the GPD parameters to depend on predictors via quantile gradient boosting. The approach is “validated” by comparing the gbex results with classical methods, such as quantile linear regression (Koenker, 2005), showing superior calibration for the extremes.
There is much to admire in Velthoen et al. (2023), but the analyses to follow depart in several important respects. The setting is above the Arctic Circle, where climate and weather processes differ substantially. Consequently, extreme events are not the primary focus; quantile gradient boosting is employed for other purposes.
No enhancements from parametric statistical procedures are used. As Breiman (2001) emphasizes, there is considerable skepticism about statistical modeling unless it has been thoroughly vetted. Nor is algorithmic tuning employed. Therefore, post–model-selection bias is avoided (Kuchibhotla, Kolassa and Kuffner, 2022).
Time-series data commonly exhibit substantial temporal dependence. To obtain valid assessments of forecasting accuracy and uncertainty, the training procedure is expanded to account for this dependence. The analyses adopt standard time-series methods (Box et al., 2015) as a training extension.
Finally, valid estimates of forecasting uncertainty are obtained using adaptive conformal prediction regions (Romano, Patterson and Candès, 2019). The requisite exchangeability is addressed explicitly. Supplemental results are provided that may be of particular interest to stakeholders.
4.2 Some Details
The Arctic multiple time-series data described above are analyzed using quantile gradient boosting, with the 60th percentile (i.e., $Q(0.60)$) as the estimation target. Let $y$ denote the numeric response variable, $\hat{y}$ its fitted value, and $\tau$ the target conditional quantile. Quantile gradient boosting minimizes the following loss function (Koenker and Bassett Jr, 1978):
[…]
With $\tau=0.60$, underestimates receive $0.6/0.4=1.5$ times more weight than overestimates when the loss is computed. For this application, about 40% of the 2023 Longyearbyen temperatures exceed $0^{\circ}\mathrm{C}$. Because relatively high values tend to be underestimated and relatively low values overestimated, fitting the conditional 0.60 quantile forces the gradient-boosting algorithm to work harder to avoid underestimates. This occurs most often among the warmest 40% of temperatures. Indirectly, therefore, these “melting days” are weighted more heavily than the rest. Melting days are a particular concern because practical climate adaptations at a local level are often needed.
If melting days are so important, one might argue for reformulating the analysis as a classification problem—above or below $0^{\circ}\mathrm{C}$. While melting is indeed a critical threshold, the degree to which that threshold is exceeded also matters. The melting process is highly nonlinear and difficult to characterize, but current evidence suggests that melting can increase at an increasing rate as temperatures rise (Polashenski, Perovich and Courville, 2012; Pizner et al., 2024).
4.3 Some Details
The Arctic multiple time-series data described above are analyzed using quantile gradient boosting, with the 60th percentile (i.e., $Q(0.60)$) as the estimation target. Let $y$ denote the numeric response variable, $\hat{y}$ its fitted value, and $\tau$ the target conditional quantile. Quantile gradient boosting minimizes the following loss function (Koenker and Bassett Jr, 1978):
[…]
With $\tau=0.60$, underestimates receive $0.6/0.4=1.5$ times more weight than overestimates when the loss is computed. For this application, about 40% of the 2023 Longyearbyen temperatures exceed $0^{\circ}\mathrm{C}$. Because relatively high values tend to be underestimated and relatively low values overestimated, fitting the conditional 0.60 quantile forces the gradient-boosting algorithm to work harder to avoid underestimates. This occurs most often among the warmest 40% of temperatures. Indirectly, therefore, these “melting days” are weighted more heavily than the rest. Melting days are a particular concern because practical climate adaptations at a local level are often needed.
If melting days are so important, one might argue for reformulating the analysis as a classification problem—above or below $0^{\circ}\mathrm{C}$. While melting is indeed a critical threshold, the degree to which that threshold is exceeded also matters. The melting process is highly nonlinear and difficult to characterize, but current evidence suggests that melting can increase at an increasing rate as temperatures rise (Polashenski, Perovich and Courville, 2012; Pizner et al., 2024).
5 Results
The results begin with simple univariate plots to provide context for the more involved analyses. Because of the months-long alternation between daylight and darkness, Arctic temperatures fluctuate differently from those at lower latitudes. In general, the figures that follow should be largely self-explanatory.
5.1 Univariate Plots
A time series of the 2 p.m. temperatures for 2023 should reflect the expected Arctic seasonal swings. Figure 1 (left panel) shows precisely that. The light-blue irregular line interpolates the daily temperatures, while the smooth black line represents a loess smoother applied to those values. The red horizontal line marks the melting point at $0^{\circ}\mathrm{C}$. Temperatures during the nocturnal months are, on average, colder than during the diurnal months, and transitions between them are gradual. Melting temperatures are common from early June through late September.
Figure 1: 2 p.m. 2023 temperatures with the time-series plot in the left panel and the histogram with a GEV distribution overlay in the right panel.
The right panel shows a histogram of the 2 p.m. temperatures for 2023 with a generalized extreme value (GEV) distribution overlay. Temperatures range from slightly below $20^{\circ}\mathrm{C}$ to slightly above $10^{\circ}\mathrm{C}$, revealing substantial variability. The GEV overlay adds little interpretive value. The histogram is roughly symmetric, although the left tail is somewhat longer than the right. No clear outliers are apparent.
5.2 Time-Series Plot for Svalbard Temperatures in 2022, 2023, and 2024
Recall the premise that each time series of daily 2 p.m. temperatures for 2022, 2023, and 2024 represents realizations from the same underlying joint probability distribution; the underlying physics for corresponding months in those three years should be comparable. However, there could be aberrations caused by Arctic amplification.
Figure 2 provides a visual comparison of the three temperature time series. The jagged lines show the 2 p.m. daily temperatures plotted against day of year. The dashed horizontal line marks the melting temperature at $0^{\circ}\mathrm{C}$.
Figure 2: Daily 2 p.m. temperatures for 2022, 2023, and 2024 using a color palette that is color-blind friendly.
Figure 2 shows substantial overlap across the three years. Seasonal trends are captured similarly, both in temperature values and in timing. The time series move largely in lockstep from early to late summer, during which melting temperatures are nearly universal. There appears to be no compelling reason to reject the claim that each series consists of random variables drawn from the same joint probability distribution. In practice, the 2022 and 2024 temperatures serve as promising holdout samples for evaluating forecasts of the 2023 temperatures.
5.3 The Quantile GBM Fit of Temperature
Some changes in the predictor variables were required. Visibility and atmospheric pressure were dropped because of excessive amounts of missing data. Wind direction (reported in degrees) was transformed into its sine and cosine components after converting degrees to radians. This representation captures the circular nature of wind direction, for which 0° and 360° denote the same physical direction. [^5] Wind direction was recorded in degrees but is intrinsically circular, with $0^{\circ}$ and $360^{\circ}$ representing the same direction. Each direction $\theta$ was therefore converted to radians and encoded as $\sin(\theta)$ and $\cos(\theta)$. Using both components is necessary because neither $\sin(\theta)$ nor $\cos(\theta)$ alone uniquely identifies a direction (e.g., $\sin(30^{\circ})=\sin(150^{\circ})$), whereas the pair $(\cos\theta,\sin\theta)$ provides a unique and smooth representation of all possible directions on the unit circle.
With the quantile parameter fixed at $\tau=0.60$, the quantile gradient-boosting procedure in R ran efficiently. The shrinkage value was set to 0.0001, the interaction depth to 6, and the minimum number of observations in a terminal node to 6. These values were chosen to foster slow convergence so that the right tail of the temperature distribution would be fitted more accurately. Using 2022 as a holdout sample to protect against overfitting indicated that approximately 27,000 iterations were appropriate for these data. [^6] With a substantially larger shrinkage value, many fewer iterations might have been adequate. However, the best shrinkage value could not be known in advance, and empirical tuning encourages “cherry picking” that can undermine later statistical inference (Kuchibhotla, Kolassa and Kuffner, 2022).
Figure 3: Observed versus fitted 2 p.m. temperatures. Black dots are the data, the blue solid line is a loess smooth serving as a visual aid, and the red horizontal line marks the melting point at $0^{\circ}\mathrm{C}$.
Figure 3 plots the observed 2 p.m. temperatures against their fitted values. The relationship is roughly linear and positive; observed and fitted values tend to increase together. Clustering around the smoothed fitted line is somewhat tighter when the fitted values exceed the melting point, with residuals rarely larger than about $\pm 2^{\circ}\mathrm{C}$. This results in part from the larger cost for underfitting imposed by the target quantile $\tau=0.60$. There is greater sparsity on either side of the temperature range between approximately $5^{\circ}\mathrm{C}$ and $0^{\circ}\mathrm{C}$. In short, the fit quality produced by quantile gradient boosting is encouraging. [^7] Koenker and Machado (1999) discuss measures of fit for conditional-quantile models derived from the quantile loss function. However, apparently no analogue of the multiple correlation coefficient exists for quantile regression. A suitable measure of fit is introduced later, when conformal prediction regions are presented.
Two points merit emphasis. First, the apparent linearity in Figure 3 says nothing about whether the lagged predictors themselves relate linearly to temperature. In Figure 1 left panel, the $x$-axis units are days; in Figure 3, they are degrees Celsius. The former shows day-to-day variation in observed temperatures, whereas the latter shows how observed temperatures vary with their fitted counterparts. Second, lagging the predictors means they generate fitted values two weeks before the future temperatures are realized. This is not yet true forecasting, because no new unlabeled cases are involved, but it is an important step in that direction. [^8] Fitted values are sometimes called predicted values, which can cause confusion. In this paper, fitted values refer to outcomes within the training or calibration data, whereas forecasted values refer to predictions for new, unlabeled cases.
Figure 4: Observed and Fitted 2 p.m. temperatures from quantile gradient boosting plotted against day of year (2023). Black dots show the observed temperatures over time, the jagged blue line is an interpolation of the fitted values, and the solid red line marks the melting point at $0^{\circ}\mathrm{C}$.
Figure 4 provides a complementary perspective that connects more directly to the Arctic-amplification context. It reproduces the observed temperatures and fitted values from Figure 3 as time series data, plotted against day of the year. As before, black dots denote the observed 2 p.m. temperatures, the jagged blue line in an interpolation of the fitted values, and the horizontal red line marks the melting point.
Overall, the fitted values track the observed temperatures rather closely and capture temperatures above $0^{\circ}\mathrm{C}$ especially well. The unusually warm days clustering in mid-July—around $10^{\circ}\mathrm{C}$—are of particular concern to climate scientists (Semenov, 2021), and they too are fitted accurately, albeit with slight underestimation. Two weeks in advance, the fitted algorithm anticipates melting temperatures effectively, including those mid-July extremes. The fit might improve further if a larger value of $\tau$ were used, but higher quantiles (e.g., $Q(0.90)$) would rely on far sparser data, potentially leading to instability. As with Figure 3, however, true forecasting remains to be performed.
5.4 Predictor Impacts on the Fitted Values
Like all algorithms, quantile gradient boosting is not a model in the traditional statistical sense (Breiman, 2001; Kearns and Roth, 2019). Nevertheless, useful insights about associations in the data can be obtained from variable-importance plots and partial-dependence plots (Friedman, 2002).
Figure 5: Relative contribution of each lagged predictor to the fitted 2 p.m. temperatures, computed as the standardized reduction in loss attributable to each predictor.
From the variable-importance plot in Figure 5, the day counter and the 2 p.m. temperature lagged by 14 days have by far the strongest associations with the fit of subsequent 2 p.m. temperatures. Rounding slightly, the former accounts for about 40% of the total reduction in loss, and the latter for a bit more than 45%. The remaining lagged predictors together contribute a little over 10%. The sum of all contributions is approximately 100%.
The prominence of the day counter and the 14-day temperature lag is unsurprising. Seasonal patterns are captured by the counter, and the temperature’s gradual evolution over days makes its own lag a strong predictor. The other variables may still capture small or localized temporal effects that, while contributing less to fitting performance, are nonetheless systematically related to the melting temperatures.
5.5 Functional Forms
The quantile gradient-boosting algorithm learns associations between each predictor and the response, including the shapes of those relationships. The partial-dependence plots in Figure 6 show the relationships between the day counter in the left and the 14-day lagged temperature on the right with the 2 p.m. temperature response, each computed with all other predictors fixed at their means. These are the two variables that dominate the fit. The black dots represent the partial-dependence values, and the solid blue lines show a loess smooth as a visual aid.
Figure 6: Partial-dependence plots for the day counter and the 14-day lagged 2 p.m. temperature. The black dots are the partial-dependence values, and the solid blue line is a loess smooth serving as a visual aid. Other predictors are fixed at their means.
Consistent with Figure 5, strong relationships are evident. A look again at Figure 4 provides context. The counter-based fitted values range from approximately $3^{\circ}\mathrm{C}$ to $3^{\circ}\mathrm{C}$, and the lagged-temperature fitted values from about $2.5^{\circ}\mathrm{C}$ to $5^{\circ}\mathrm{C}$. The first plot is roughly symmetric and concave, peaking during the summer months, consistent with the longitudinal pattern shown in Figure 4. The second plot is S-shaped, with its steepest slope beginning immediately after the melting point at $0^{\circ}\mathrm{C}$ and plateauing near $2.5^{\circ}\mathrm{C}$ at the left and $5^{\circ}\mathrm{C}$ at the right. The lagged observed temperatures have their strongest association with fitted temperatures just above the melting point. The practical significance of this threshold will be discussed in greater depth below. [^9] The short horizontal strings of dots visible in the plot represent rounding artifacts; the values differ slightly. The remaining four predictors also exhibit nonlinear relationships with the temperature response, but because of their comparatively small contributions to the fit, their partial-dependence plots are omitted in the interest of space.
6 Forecasts and Estimated Uncertainty
Addressing uncertainty in a principled way has been anticipated in the research design and algorithmic procedures employed. The 2024 data serve as a pristine holdout sample of test data that can represent unlabeled cases used in genuine forecasting. [^10] The 2024 data have labels that are useful for didactic purposes and for evaluation of forecasting accuracy. But those labels have no role in the forecasts themselves and would not be known in an operational setting. Forecasts for the 2024 data can be produced using the previously fitted quantile gradient-boosting algorithm and its predict() function in R. From the perspective of the trained algorithm, the 2024 lagged predictors constitute new, unlabeled data.
Obtaining the adaptive conformal prediction regions is more involved. All conformal prediction regions begin with a pre-specified coverage probability. By convention, this probability is denoted by $(1-\alpha)$, where $\alpha$ lies between 0 and 0.50. In practice, coverage probabilities range from slightly above 0.50 to nearly 1.0, with values between 0.75 and 0.95 most common. The value of $\alpha$ is determined by subjective considerations shaped by the data and the application. For the conformal approach used with the Svalbard data, the coverage is $1-\alpha=0.80$.
Larger coverage probabilities lead to larger conformal prediction regions. If the goal is to be highly confident that a forecast will be found within a particular prediction region, it makes sense that the length of the prediction region should be long. If the goal is to reduce the length of the prediction region in service of greater precision, confidence that the forecast will be found within that region must decline. Conformal prediction regions formalize this trade-off between coverage and precision. Stakeholders might legitimately prefer a different value for coverage, and within limits imposed by the data, another coverage value is easily implemented. [^11] Ideally, a future decision about a coverage probability would be made with new data. If a new coverage probability is selected for the 2024 data after seeing the results, new complications are introduced and the conformal methods would need some modest changes (Sarkar and Kuchibhotla, 2023).
Conformal prediction regions are constructed from “nonconformal scores.” These scores can be derived directly from the calibration-data residuals when the calibration data are used as the input into the previously trained boosting algorithm (i.e., predict() in R). Because temperatures from the Svalbard calibration data have strong temporal dependence and the boosting residuals do as well, the residuals were whitened using a first-order autoregressive model (AR(1)). This removes the short-range serial dependence while preserving the level and variability required for valid nonconformal scores. [^12] These steps were guided by estimated autocorrelation functions and Box–Ljung tests determining whether the AR(1) model produced white noise residuals.
Exchangeability is required for valid prediction regions (Vovk, Gammerman and Shafer, 2005; Vovk et al., 2017; Angelopoulos, Barber and Bates, 2024). One implication of exchangeability is that it does not matter, for the theoretical guarantees, that a test observation is realized after all of the calibration observations. A formal summary of conformal prediction is provided in Appendix A along with some computation details. Angelopoulos and Bates (2023) provide an accessible introduction to conformal prediction.
Figure 7 illustrates how conformal prediction regions work. The figure highlights important features of the Svalbard data but is not meant to be a reproduction of those data. On the horizontal axis is the forecasted temperature. On the vertical axis is the observed future temperature that would, in practice, be unknown when a forecast is made. The solid blue line conveys how the forecasts and future temperatures are systematically related. The dashed red lines are the upper and lower bounds of the adaptive prediction region for each forecast. If $1-\alpha=0.80$, the prediction region for any forecast contains the correct future temperature with a probability of at least 0.80.
Figure 7: Simulation of adaptive conformal prediction regions from quantile gradient boosting with $\tau=0.80$. The horizontal axis is the forecasted temperature. The vertical axis is the observed future temperature. The solid blue line shows how forecasts and true future temperatures are systematically related. The dashed red lines show the upper and lower bounds of the adaptive prediction region, and the two vertical dotted magenta lines mark the prediction regions when the forecasted temperature happens to be $10^{\circ}\mathrm{C}$ or $5^{\circ}\mathrm{C}$.
To help fix these ideas, consider two illustrative forecasts, one at $10^{\circ}\mathrm{C}$ and the other at $5^{\circ}\mathrm{C}$. Several features of their prediction regions are apparent and apply to all forecasted temperatures in the figure. Forecasts are not necessarily in the middle of their prediction regions, which means that the regions need not be symmetric around the forecast. In addition, because for the quantile gradient-boosting algorithm $\tau$ was set to 0.60, the forecasts tend to fall above each region’s median. Finally, the lengths of the adaptive prediction regions grow shorter as the fit improves (Romano, Patterson and Candès, 2019). In Figure 7, forecasts above $0^{\circ}\mathrm{C}$ fit the observed temperatures better and have shorter region lengths. In the Svalbard setting, forecasts would be produced one day at a time, but Figure 7 is meant to convey the results after many forecasts are made. The figure also makes plain that there is no requirement that the forecasts be correct.
Although the nominal tail probabilities defining the lower and upper limits are symmetric (e.g., $\alpha/2$ and $1-\alpha/2$), the resulting prediction regions need not be. The asymmetry arises because the values of the upper and lower bounds are determined by the empirical distribution of the nonconformal scores themselves. Their ranks determine the quantiles, but the nonconformal scores need not be equally spaced from one another. For example, if the large positive scores are less dispersed than the large negative scores, the prediction region above the forecasts will be shorter than the prediction region below the forecast.
6.1 Uncertainty Results
The lengths of the 2024 conformal prediction regions vary substantially, consistent with Figure 3. For days in which the forecasted 2 p.m. temperature exceeds 0, the average half-width is about $\pm 2.6^{\circ}\mathrm{C}$, and the half-widths range from about $1.4^{\circ}\mathrm{C}$ to $6.2^{\circ}\mathrm{C}$. For days in which the forecasted 2 p.m. temperature does not exceed 0, the average half-width is about $\pm 4.7^{\circ}\mathrm{C}$, and the half-widths range from about $2.0^{\circ}\mathrm{C}$ to $8.3^{\circ}\mathrm{C}$. Performance is better for the warmer diurnal temperatures because $\tau>0.50$.
Stakeholders might not find much comfort in these summary statistics. Even if the mean half-widths promise some guidance for policy and practice, the ranges of the half-widths counsel caution. Perhaps more useful information could be found in estimates of the distributions of the future observed temperatures for each forecasted temperature. In principle, there would be more information, but the sample sizes would be too small, even over several years of data, to support such estimates properly.
In the spirit of conditional distributions of future temperatures, a more modest approach might work. The forecasts can be binned to increase the sample sizes, and rather than estimating full conditional distributions, estimate how likely a future temperature of $0^{\circ}\mathrm{C}$ is exceeded. Treating melting as a “positive”, when a forecasted temperature is above $0^{\circ}\mathrm{C}$ and the future temperature $0^{\circ}\mathrm{C}$, a true positive results. When the forecasted temperature is $\leq 0^{\circ}\mathrm{C}$ and the future temperature is $0^{\circ}\mathrm{C}$, one has a false negative. Ideally, true positives are far more common than false negatives.
Figure 8: For the 2024 holdout sample, the mean forecasted temperature in a bin is on the horizontal axis. The percentage of cases in each of 20 equal-width bins having a true future temperature exceeding $0^{\circ}\mathrm{C}$ is on the vertical axis. Filled circles correspond to bins with mean forecasted temperature above $0^{\circ}\mathrm{C}$. Open circles correspond to those below. The solid blue line is a loess smoother.
Figure 8 provides a visual summary of these ideas using the 2024 holdout test data. The horizontal axis shows the mean of the daily forecasted temperatures for each of 20 equal-width bins. Because the future temperatures are known for the 2024 holdout cases, the vertical axis shows the percentage of forecasts within each bin whose true future temperature exceeds $0^{\circ}\mathrm{C}$. (The 14-day lag of each predictor ensures that these are future outcomes.) The loess overlay is S-shaped, with the greatest rate of change between about $2^{\circ}\mathrm{C}$ and $2^{\circ}\mathrm{C}$.
The conclusions from Figure 8 are straightforward. The figure is a confirmation of the analyses above because there is strong evidence of forecasting skill in the holdout sample. As the forecasted temperature increases from about $10^{\circ}\mathrm{C}$ to about $10^{\circ}\mathrm{C}$, the percentage of cases above the true future melting temperature increases in a nearly monotonic manner from 0.0% to 100%.
Moreover, the percentages on the vertical axis in combination with the smoother confirm that true positives are far more common than false negatives. The largest percentage of false negatives is concentrated between about $2^{\circ}\mathrm{C}$ and $1^{\circ}\mathrm{C}$ where a little less than 20% of the future true temperature values exceed the melting point. As soon as the forecasted temperature exceeds $0^{\circ}\mathrm{C}$, more than 80% of the forecasts in a bin correspond to future temperatures above $0^{\circ}\mathrm{C}$, reaching 100% as the bin mean exceeds $4^{\circ}\mathrm{C}$. The large vertical gap between filled and open circles highlights the relatively abrupt phase transition centered on $0^{\circ}\mathrm{C}$ responsible for the dramatic increases in true positives. Because melting is such an important event, stakeholders might find Figure 8 encouraging. Some kinds of temperature forecasts might be accurate enough to be useful.
But physical interpretations can be subtle. When sea ice forms, its hexagonal lattice cannot readily incorporate salt ions. The excluded salt forms brine pockets and channels among the ice grains, leaving the ice as nearly fresh water while the brine becomes increasingly salty. As freezing continues, the concentrated brine’s freezing point can fall well below $0^{\circ}\mathrm{C}$. Before that stage, a two-phase mixture of nearly pure ice and brine remains. Melting reverses this process: pure ice begins melting at $0^{\circ}\mathrm{C}$, whereas frozen brine melts at lower temperatures.
The rates of melting and freezing depend on cloud cover, winds, and ocean currents that transport warmer or cooler water to local sites. Consequently, freezing and melting occur over a range of temperatures and time scales rather than at a fixed threshold. This interpretation aligns with Figure 8.
From a policy perspective, even approximate information about when melting will occur can be valuable. Although graphs like Figure 8 cannot be produced prospectively—since future temperatures are unknown—they can be constructed retrospectively using historical weather-station data. Such analyses help stakeholders gauge likely rates of true positives and false negatives conditional on forecasted temperature. For example, if a forecast exceeds $4^{\circ}\mathrm{C}$, the realized future temperature will almost certainly be $0^{\circ}\mathrm{C}$. [^13] The plotted percentages should not be interpreted as probabilities. Temporal dependence persists among the forecasted temperatures within each bin, although its structure is not easily characterized. The forecasts are ordered by temperature rather than by time, and the time gaps between forecasts within a bin can range from days to months. For example, forecasted temperatures during late spring and early fall may be similar even though they occur several months apart.
Given these results, the role of conformal prediction regions can be briefly revisited. For a new, unlabeled case T+1T{+}1, the true temperature will fall within its adaptive conformal prediction region with probability at least $1-\alpha$. When the lower bound of a prediction region is greater than zero, the probability that the future temperature will be above 0 is $1-\alpha/2$. For a specified coverage probability of 0.80, that probability would increase to 0.90. Coverage is gained because under some circumstances, the prediction region’s upper bound may not matter much. This approach could work well for Svalbard insofar as temperatures $0^{\circ}\mathrm{C}$ are far more concerning than lower temperatures. In the spirit of Figure 8, when the lower bound of the prediction region is greater than $0^{\circ}\mathrm{C}$, one may simply assume that the forecast of melting is correct.
7 Discussion
The analysis of 2 p.m. temperatures in Svalbard, Norway, combines three interlocking statistical traditions: gradient boosting with a quantile as the estimation target, an AR(1) model to construct exchangeable nonconformal scores, and adaptive conformal prediction regions computed with quantile random forests, which is robust to sparse data. To the best of the author’s knowledge, the combined use of three temporally separated samples, quantile gradient boosting, AR(1)-based whitening, and regime-specific adaptive conformal prediction has not previously appeared in the literature. The statistical procedures were informed by extensive knowledge of the research site. Equally important was guidance from climate science in the choice of algorithms and in interpreting their output. The aim was to contribute both to scientific understanding and to practical human adaptations to Arctic melting.
A two-week forecasting lead time before the onset of widespread surface melting is a very short fuse, but it would give Arctic community administrators critical time to prepare for infrastructure resilience, public safety, and logistical challenges (Streletskiy, Shiklomanov and Christiansen, 2019; U.S. Arctic Research Commission, 2003). Local governments could limit heavy-vehicle use on thaw-sensitive roads, runways, and around pipelines; pre-position maintenance materials; and inspect culverts, bridges, and drainage channels before damage occurs (Arctic Research Consortium of the United States (ARCUS), 2021). Water managers could adjust reservoir levels or activate temporary treatment measures to mitigate siltation and contamination from meltwater inflows.
In communities dependent on ice roads, snowmobile trails, or frozen river crossings, early warnings would permit orderly resupply and fuel delivery before surface travel becomes unsafe (Chen et al., 2025). Coastal and hillside settlements could prepare for increased risks of shoreline erosion or permafrost-related landslides (The Climate Institute and The Firelight Group, 2022). Health authorities might use the lead time to issue advisories on water quality or vector-borne disease risk as standing water accumulates. Because of Arctic warming, even Iceland is now reported to host mosquitoes (Straker, 2025).
More broadly, forecasts on a two-week timescale could foster coordination among local administrators, regional governments, and research stations, improving situational awareness and resource allocation across sectors. Even modest improvements in lead time may yield substantial adaptive benefits in Arctic regions where melt conditions evolve rapidly and logistical flexibility is limited. [^14] The last three paragraphs were informed by a literature search undertaken by ChatGPT, which also provided initial wording later edited for consistency and style.
8 Conclusions
Projections of future global warming have long been central to IPCC assessments, complemented by many smaller-scale forecasting efforts. The analyses presented here contribute to the latter. Temperature forecasts with a two-week lead time appear promising, especially when focused on Arctic thresholds associated with widespread melting. Forecasting uncertainty is explicitly addressed, and results are presented in a manner intended to be accessible and useful to stakeholders.
The analyses also illustrate that one need not rely on “industrial-strength” data or computational infrastructure to obtain meaningful forecasts. With thoughtfully applied algorithms and accessible data, informative short-term forecasts can be generated on an ordinary desktop computer with an internet connection.
Appendix A. Statistical Narrative
The adaptive conformal forecasting procedure proceeds sequentially in the following steps. Separate calibration and forecasting steps ensure that all prediction intervals are based solely on information available at the time of forecasting. The sequence below is written with R in mind but also applies to Python.
  1. 1.
    Data preparation. Subset the longitudinal Longyearbyen weather-station data into training data from 2023, calibration data from 2022, and test data from 2024. Specify a coverage probability $1-\alpha$. For each year, the response $y_{t}$ is the daily 2 p.m. air temperature in degrees Celsius. The predictors for each dataset $X_{t-14}$ are the same meteorological variables lagged by 14 days.
  2. 2.
    Algorithm training. With $\tau=0.60$, fit quantile gradient boosting to the 2023 training data (train), yielding a trained algorithm (gbm) and an optimal iteration count (best.iter).
  3. 3.
    Residuals for calibration. Using the 2022 calibration data (calibration), compute fitted values
    $\widehat{y}_{t}^{2022}=\texttt{predict}(\texttt{gbm},\texttt{calibration},\texttt{best.iter}),$
    and residuals
    $r_{t}^{2022}=y_{t}^{2022}-\widehat{y}_{t}^{2022}.$
  4. 4.
    Regime split. In response to subject-matter concerns, split the calibration residuals into warm and cool regimes using the fitted values:
    $\text{Warm if }\widehat{y}_{t}^{2022}>0,\qquad\text{Cool if }\widehat{y}_{t}^{2022}\leq 0.$
  5. 5.
    Temporal-dependence adjustment. Because of temporal dependence in $r_{t}^{2022}$, fit a first-order autoregressive model with intercept to the full residual sequence,
    $r_{t}^{2022}=\alpha+\phi r_{t-1}^{2022}+\varepsilon_{t}.$
    This produces a single set of AR(1) innovations. These innovations are then allocated to the warm and cool regimes according to the fitted-values split, which removes short-range serial dependence without altering the level or variability.
  6. 6.
    Innovations (nonconformal scores). Extract the AR(1) innovations $\varepsilon_{t}$ for each regime. These approximately white-noise innovations serve as the nonconformal scores.
  7. 7.
    Quantile-function estimation. For each regime, fit a quantile random forest (QRF) to the 2022 nonconformal scores, using the 2022 fitted values and the 2022 lagged predictors as covariates. QRF learns how the variability of the nonconformal scores depends on the fitted temperature level and on the lagged meteorological predictors.
  8. 8.
    Forecast generation. Using the 2024 test data (forecast), compute gbm forecasts
    $\widehat{y}_{t}^{2024}=\texttt{predict}(\texttt{gbm},\texttt{forecast},\texttt{best.iter}),$
    and assign each case to the warm or cool regime according to the sign of its forecast.
  9. 9.
    Nonconformal-score quantiles. For each 2024 case, supply its forecasted value and lagged predictors to the corresponding regime’s QRF to obtain predicted score quantiles $\widehat{q}_{\alpha/2}$ and $\widehat{q}_{1-\alpha/2}$.
  10. 10.
    Prediction intervals. Construct lower and upper adaptive conformal prediction bounds:
    $\mathrm{PI}_{\mathrm{lower}}=\widehat{y}_{t}^{2024}+\widehat{q}_{\alpha/2},\qquad\mathrm{PI}_{\mathrm{upper}}=\widehat{y}_{t}^{2024}+\widehat{q}_{1-\alpha/2}.$
    The interval widths reflect both the variability captured by QRF and any remaining dispersion in the nonconformal scores.
These linked procedures combine quantile gradient boosting, AR(1) whitening, and quantile-random-forest calibration to yield adaptive conformal prediction regions. The lower and upper bounds use the ranks of the nonconformal scores, while the magnitudes of those scores determine the lengths of the resulting prediction regions.
Exchangeability of Nonconformal Scores
The supervised learning algorithm is first fitted on the training data $\{(X_{t},Y_{t})\}_{t=1}^{T_{\mathrm{train}}}$. Predictions are then generated for the separate calibration data $\{X_{t}\}_{t=1}^{T}$, and the resulting residuals or their AR(1) innovations form the nonconformal scores $\{S_{t}\}_{t=1}^{T}$.
If $\{S_{t}\}_{t=1}^{T+1}$ are (approximately) exchangeable—perhaps after whitening— then, conditional on the fitted algorithm, the calibration scores and the new score $S_{T+1}$ are also exchangeable. Consequently,
$\Pr\!\left\{\,S_{T+1}\leq q_{1-\alpha}\big(\{S_{t}\}_{t=1}^{T}\cup\{S_{T+1}\}\big)\right\}\geq 1-\alpha,$
and the adaptive conformal prediction region attains marginal coverage of at least $1-\alpha$.
Exchangeability means that the joint distribution of the nonconformal scores is invariant to permutations of their indices. Although in practice the scores are realized in temporal order, their joint law would be unchanged had they been realized in any other order. That a new unlabeled case is realized at time T+1T{+}1 is immaterial. For the analyses of the Svalbard data, exchangeability depends on the whitening of the calibration data residuals. The residuals from the training data are not exchangeable, and neither are the residuals from the calibration data before whitening.